OpenStudy (anonymous):

log(3x+5)-log(x-5)=log(8)

OpenStudy (anonymous):

i got it so log(3x+5/x-5)=log8 and dont know where to go from there

Nnesha (nnesha):

hint: \[\huge\rm log_b x = \log_b y\] \[\huge\rm\cancel { log_b} x = \cancel{\log_b} y\] x=y

OpenStudy (anonymous):

so 3x-5/x-5=8

Nnesha (nnesha):

yep now simple algebra

OpenStudy (anonymous):

i got -20 but the answer is 9

Nnesha (nnesha):

how did you get -20?

OpenStudy (anonymous):

i cross multiplied

OpenStudy (anonymous):

so 8x-40=3x+5

Nnesha (nnesha):

solve for x

OpenStudy (anonymous):

i got -20

Nnesha (nnesha):

show ur work

OpenStudy (anonymous):

oh i found my mistake thank you!

Nnesha (nnesha):

r u sure ? u got it ?

OpenStudy (anonymous):

yep! can i ask you one more though?

Nnesha (nnesha):

alright

OpenStudy (anonymous):

log8+3logx=3 which i know is log8+logx^3=3 but what do you do if theres no log on the other side?

Nnesha (nnesha):

apply log properties which one you should apply ?

OpenStudy (anonymous):

change of base?

Nnesha (nnesha):

yeah or convert log to exponential form just like we did before

Nnesha (nnesha):

nope there is a variable so you can't use change of base formula

OpenStudy (anonymous):

how would you do that though because isnt it 8x^3=3

Nnesha (nnesha):

log (8x^3) = 3

Nnesha (nnesha):

\[\log = \log_{10}\]

OpenStudy (anonymous):

but idk how you would make that exponential

Nnesha (nnesha):